OpenStudy (kiki99):

After years of overhunting, environmental scientists have reintroduced bears to Yellowstone National Park. The initial number of bears reintroduced to the park was 20 and, after 10 years, the population is estimated to be around 110. Part A: Assuming an exponential growth pattern, use the information given to write the model for this situation.

OpenStudy (aveline):

Let's use \[y=A(1+r)^ {t}\] where A represents the initial population, r represents the rate of growth, t represents the time interval, and A represents the total population. The initial number, A, is 20. The time interval, t, is 10 years. The total population, y, is 110. So now we can fill in the equation: \[110=20(1+r)^ {10}\]

OpenStudy (kiki99):

okay thanks i was just about to put my result in and ask if it was right

OpenStudy (kiki99):

can you help me with part B? @Aveline

OpenStudy (kiki99):

What is the annual growth rate (rounded to the nearest tenth of a percent) of the new bear population in Yellowstone National Park?

OpenStudy (aveline):

Yeah.

OpenStudy (aveline):

We have our equation \[110=20(1+r) ^{10}\] First, we divide 20 from each side: \[5.5=(1+r)^ {10}\] Can you simplify from here?

OpenStudy (kiki99):

5.5/20=110

OpenStudy (kiki99):

what do i do to the other side what do i divide by 20

OpenStudy (aveline):

I already divided 20 for you ;)

OpenStudy (kiki99):

oh okay no we could not simplify any further

OpenStudy (kiki99):

is that it? @Aveline

OpenStudy (aveline):

You need to simplify the exponents.

OpenStudy (kiki99):

@Aveline i dont remember how